 Use inductive reasoning to make a conjecture about a rule that relates the number you selected to final answer. Try to prove your conjecture by using deductive reasoning.


1. Pick a number


Double it


Subtract 20 from the answer


Divide by 2


Subtract the original number Result:


2. Pick a number Multiply it by 9


Add 21


Divide by 3


Subtract three times the original number Result:


3. Pick a number


Add 6


Multiply the answer by 9


Divide the answer by 3


Subtract 3 times the original number Result




1
Expert's answer
2022-01-06T12:02:34-0500

1.

Let the number be "1."

Double it

"2\cdot1=2"

Subtract 20 from the answer


"2-20=-18"

Divide by 2


"\dfrac{-18}{2}=-9"

Subtract the original number


"-9-1=-10"

Result:

"-10"

Let the number be "-\dfrac{1}{3}."

Double it

"2\cdot(-\dfrac{1}{3})=-\dfrac{2}{3}"

Subtract 20 from the answer


"-\dfrac{2}{3}-20=-\dfrac{62}{3}"

Divide by 2


"\dfrac{-\dfrac{62}{3}}{2}=-\dfrac{31}{3}"

Subtract the original number


"-\dfrac{31}{3}-(-\dfrac{1}{3})=-10"

Result:

"-10"

Let the number be "\sqrt{2}."

Double it

"2\cdot(\sqrt{2})=2\sqrt{2}"

Subtract 20 from the answer


"2\sqrt{2}-20=2\sqrt{2}-20"

Divide by 2


"\dfrac{2\sqrt{2}-20}{2}=\sqrt{2}-10"

Subtract the original number


"\sqrt{2}-10-\sqrt{2}=-10"

Result:

"-10"

A conjecture: the result is always "-10."


Take arbitrary number "c\in\R."

Double it

"2\cdot(c)=2c"

Subtract 20 from the answer


"2c-20=2c-20"

Divide by 2


"\dfrac{2c-20}{2}=c-10"

Subtract the original number


"c-10-c=-10"

Result:

"-10"


We see that the result does not depend on the taken number.

Therefore we proved that the result is always "-10" regardless of the number we pick.


2.

Let the number be "1."

Multiply it by 9

"1\cdot9=9"

Add 21


"9+21=30"

Divide by 3


"\dfrac{30}{3}=10"

Subtract three times the original number


"10-3(1)=7"

Result:


"7"

Let the number be "\dfrac{1}{2}."

Multiply it by 9

"\dfrac{1}{2}\cdot9=\dfrac{9}{2}"

Add 21


"\dfrac{9}{2}+21=\dfrac{51}{2}"

Divide by 3


"\dfrac{\dfrac{51}{2}}{3}=\dfrac{17}{2}"

Subtract three times the original number


"\dfrac{17}{2}-3(\dfrac{1}{2})=7"

Result:


"7"

Let the number be "\sqrt{2}."

Multiply it by 9

"\sqrt{2}\cdot9=9\sqrt{2}"

Add 21


"9\sqrt{2}+21=9\sqrt{2}+21"

Divide by 3


"\dfrac{9\sqrt{2}+21}{3}=3\sqrt{2}+7"

Subtract three times the original number


"3\sqrt{2}+7-3(\sqrt{2})=7"

Result:


"7"

A conjecture: the result is always "7."


Take arbitrary number "c\in\R."

Multiply it by 9

"c\cdot9=9c"

Add 21


"9c+21=9c+21"

Divide by 3


"\dfrac{9c+21}{3}=3c+7"

Subtract three times the original number


"3c+7-3c=7"

Result:

"7"


We see that the result does not depend on the taken number.

Therefore we proved that the result is always "7" regardless of number we pick.


3.

Let the number be "1."

Add 6


"1+6=7"


Multiply the answer by 9

"7\cdot9=63"

Divide the answer by 3


"\dfrac{63}{3}=21"

Subtract 3 times the original number


"21-3(1)=18"

Result:


"18"



Let the number be "-\dfrac{1}{2}."

Add 6


"-\dfrac{1}{2}+6=\dfrac{11}{2}"


Multiply the answer by 9

"\dfrac{11}{2}\cdot9=\dfrac{99}{2}"

Divide the answer by 3


"\dfrac{\dfrac{99}{2}}{3}=\dfrac{33}{2}"

Subtract 3 times the original number


"\dfrac{33}{2}-3(-\dfrac{1}{2})=18"

Result:


"18"



Let the number is "\sqrt{2}."

Add 6


"\sqrt{2}+6=\sqrt{2}+6"


Multiply the answer by 9

"(\sqrt{2}+6)\cdot9=9\sqrt{2}+54"

Divide the answer by 3


"\dfrac{9\sqrt{2}+54}{3}=3\sqrt{2}+18"

Subtract 3 times the original number


"3\sqrt{2}+18-3(\sqrt{2})=18"

Result:


"18"


A conjecture: the result is always "18."


Take arbitrary number "c\in\R."

Add 6


"c+6=c+6"


Multiply the answer by 9

"(c+6)\cdot9=9c+54"

Divide the answer by 3


"\dfrac{9c+54}{3}=3c+18"

Subtract 3 times the original number


"3c+18-3(c)=18"

Result:


"18"


We see that the result does not depend on the taken number.

Therefore we proved that the result is always "18" regardless of number we pick.